# A Deep Dive into Large Language Models for Automated Bug Localization and Repair

### v2

@@ 3.1.2. Baseline
For our baselines, we selected the two best-performing models, based on our current knowledge. Given that the Tufano Small and Medium datasets are part of the CodeXGLUE benchmark (CodeXGLUE-Leaderboard, ), our selection was based on the highest-ranking models from the leaderboard. For the CodeReviewer dataset, our baselines include the CodeT5 model as reported in (li2022codereviewer, ) and NSEdit (hu2022fix, ) model. These models achieve the top two accuracies on the CodeReviewer dataset.
The NSEdit model (hu2022fix, ), a 12-layer encoder-decoder with 223M parameters, expands to around 500M including the re-ranking model. It achieved 23.86% EM accuracy on the Tufano Small dataset and 13.46% on the Tufano Medium. Another baseline, CoText (phan2021cotext, ), also has 223M parameters but supports up to 11B, achieving 22.64% on the small and 15.36% on the medium dataset.
On the CodeReviewer with comment dataset, the CodeT5 model achieved an accuracy of 24.42%, as reported in the original paper (li2022codereviewer, ) [^1] In the CodeReviewer paper, specialized additional pre-training increased accuracy to 30.32%. Since the pre-training was custom and the dataset is not public, we compare our results with their reported CodeT5 results.. For the CodeReviewer without comment dataset, the original paper did not report any accuracy (li2022codereviewer, ). However, replicating their method on this dataset is straightforward and yields an accuracy of 11.48%. For both CodeReviewer with and without comment datasets, we run the NSEdit model (hu2022fix, ) and report the accuracies in Table 1.
@@ 3.2.1. Experimental Setup
Dataset. Defects4J is a benchmark dataset consisting of real-world Java bugs from 17 projects. This dataset is widely used for evaluating automated program repair methods. From Defects4J, we primarily focus on ‘single-hunk’ bugs, which can be fixed with changes in a single contiguous block of code. This type of bug is well-suited for APR methods and has been extensively used in prior research (coconut, ; cure, ; dlfix, ; rewardrepair, ; knod, ). This study considers a total of 240 single-hunk bugs from the Defects4J benchmark.
Patch generation with fine-tuned model. For location prediction, we employ our fine-tuned CodeT5 bug localization model. As we did not fine-tune the model on the Defects4J dataset, to ensure diversity of sample distribution seen by the model, we fine-tune it on different datasets: Tufano Small, Tufano Medium, and GitHub (zhu2021syntax, ). For bug fixing task, we select CodeParrot, a 110M parameter model, due to its superior performance over the similarly sized CodeGPT-110M and comparable performance to larger models like CodeGen and PolyCoder. Considering the resource-intensive nature of larger models, we used the CodeParrot model fine-tuned on same three datasets.
To generate patches for Defects4J bugs, we first predict bug locations with three fine-tuned CodeT5 models. Then, similar to RQ1, we construct prompt 3 (¡truncated bug¿ [sep] ¡shared prefix¿), as shown in Figure 5. Each fine-tuned bug fixing model then generates 70 patches per input prompt, i.e., buggy sample, resulting in a total of 210 patches for each bug. This quantity is relatively small compared to the number of patches generated by most existing APR techniques (coconut, ; cure, ; rewardrepair, ; xia@alpharepair, ; knod, ; tenure, ) evaluated on Defects4J.
Accuracy computation via test execution. Unlike RQ1, where accuracy was determined via token-by-token match, this experiment utilizes the test cases from the Defects4J framework to validate the generated fixes. A patch is considered correct if it successfully passes all test cases. To further enhance our confidence, we manually executed all corrective patches to ensure they indeed pass all tests and effectively fix the bugs.
@@ 3.2.2. Results
CURE (cure, ) 18 32 37 49 52 70
RewardRepair (rewardrepair, ) 28 39 49 58 73 85
Recoder (zhu2021syntax, ) 36 51 57 64 85 -
KNOD (knod, ) 33 49 62 70 84 100
Tare (tare, ) - - - - - 109
AlphaRepair (xia@alpharepair, ) - - - - - 110
TENURE (tenure, ) - - - - - 129
@@ 3.3.1. Experimental Setup
To solely assess the impact of our four designed prompts (shown in Figure 5) on bug fixing accuracy, we utilize the ground truth locations of buggy tokens instead of predicted ones. This method ensures a controlled setup, specifically designed to prevent location prediction errors from influencing the accuracy of bug fixes. We fine-tune six bug-fixing LLMs with each of the four prompts on the Tufano Small dataset (tufano2019empirical, ) from the CodeXGLUE benchmark (lu2021codexglue, ). Finally, we report their test accuracy in Table 4.
@@ 3.3.2. Results
In Table  4, each row shows the bug-fixing accuracy of a specific LLM, and each column shows the accuracy achieved with each type of prompt. CodeGPT starts with a lowest accuracy of 16.07% using prompt 1 but shows a consistent and significant improvement across prompts, reaching 56.98% with prompt 4. CodeParrot begins at 24.5% accuracy with prompt 1, marking the highest accuracy among all models for this prompt. Accuracy for prompts 2 and 3 remains similar; however, with prompt 4, CodeParrot reaches 63.83%, the highest across all models for this prompt. CodeGen-350M exhibits a similar pattern, beginning at 22% and consistently improving to achieve up to 60.22% accuracy. PolyCoder also shows this trend, starting at 23.66% and consistently improving to achieve an accuracy of 62.29%.
@@ 4.1. Deep Learning-Based Automated Program Repair
Research in deep learning-based program repair has primarily focused on three approaches: neural machine translation (NMT)-based, edit-based, and the more recent large language model (LLM)-based methods. NMT-based approaches treat the correction of code as a translation task, where the ‘fixed’ code is translation target. This method is relatively simple to implement because both the input (buggy code) and the target (fixed code) are sequences of code, enabling direct modeling as a sequence-to-sequence (Seq2Seq) task (Chen2019sequencer, ; coconut, ; cure, ; rewardrepair, ). However, a potential downside of directly predicting the fixed code is that it may encourage the models to learn the copying behavior due to the typical large overlap between the buggy and fixed code. This copying behavior causes over-fitting and overlooks the goal of editing to change the code and fix the bug.
Editing-based approaches try to directly target the portion of the code that contains bug and usually contain an explicit mechanism for bug localization. Although edits are often shorter than the fixed code, most of the existing approaches often require multiple stages of editing with complicated models (hashimoto2018retrieve, ; zhu2021syntax, ; yin2018learning, ; tarlow2020learning, ; knod, ; tare, ). They may also rely on a graph representation of the buggy code (yao2021learning, ; Dinella2020, ), with the implicit assumption that the buggy code can be parsed into a graph, which may not be true if bugs contain syntactical errors.
Recent studies have explored large language models (LLMs) for automated program repair (APR), showing that LLMs, either as-is or with straightforward fine-tuning, have good fixing capabilities (jiang2023impact, ; xia2023automated, ). However, these studies did not address bug localization or its integration with bug fixing, nor did they conduct a comprehensive study on the design of prompts for fine-tuning LLMs.
@@ 4.2. Large Language Models For Software Engineering
LLMs have been applied to several software engineering tasks in addition to APR, including code generation (related-generation-1, ; related-generation-2, ; related-generation-3, ; related-generation-4, ), vulnerability detection and fixing (related-vul-1, ; related-vul-2, ; related-vul-3, ), fuzzing (related-fuzz, ) and software specification generation (related-specification, ). Such adoption of LLMs typically requires fine-tuning if a decent amount of data is available, or prompt design if the prompt engineering approach is utilized.
@@ 5. Threats to Validity
In our research, we have investigated the effectiveness of our bug fixing framework using several open-source, widely recognized datasets such as CodeXGLUE  (lu2021codexglue, ) and CodeReviewer  (li2022codereviewer, ). These datasets present a diverse set of bugs collected from GitHub and span various programming languages. However, there’s a possibility that our results may not generalize across other datasets. To address this threat and ensure broader applicability, we further evaluated the performance of our fine-tuned model on the unseen dataset, Defects4J, which is a widely used real-world bug dataset.
In our large-scale experiments, we have developed various tools and scripts to facilitate our experiments, and it is conceivable that they might contain bugs. To mitigate this threat, we have used widely used PyTorch and Hugging Face libraries. Additionally, we have carried out thorough validity checks and repeated each experiment several times to ensure result consistency. Furthermore, all our data and code have been made publicly accessible at https://figshare.com/s/c338829ae1b21cb9e3e3 for reviewer scrutiny.
In our study, we measured bug fix accuracy using the “exact match” criterion, a metric commonly adopted by the research community. Specifically for the Defects4J dataset, we determined accuracy through test validation. This involved executing the test cases on the generated patches; a patch was deemed correct if it passed all test cases. We are confident that this approach provides a reasonable and reliable accuracy metric.
@@ 6. Conclusion
In this study, we have explored the intricacies of LLMs in the realm of automated program repair. To sum up, this paper makes the following pivotal contributions: (1) Granularity Shift: We introduced a new direction that localizes and fixes bugs at token granularity instead of the traditional line granularity. This innovation drastically reduces the redundancy for bug-fixing models to produce shared prefixes or suffixes. (2) Prompting Strategy: We presented a novel design of four diverse prompts to optimize LLMs as bug-fixing models. This approach proves that with the right prompt design, token-granulated localization brings a potent inductive bias, leading to substantial improvements in bug-fixing accuracy. (3) Harmonizing Discrepancies: We addressed the unique challenge posed by inconsistencies between bug localization and fixing models during LLM fine-tuning. To reconcile this, we proposed an intermediate adjustment module, further strengthening the capability of our bug-fixing models. (4) In-depth Exploration: Our extensive research into LLMs’ bug-fixing capabilities encompassed experiments with four different prompts across six LLMs and five benchmarks. The culmination of this research resulted in the creation of Toggle, our Token-Granulated Bug Localization and Repair framework. Toggle has surpassed existing state-of-the-art techniques across several benchmark assessments.
Our proposed framework, Toggle demonstrates the success of our strategy, establishing a new benchmark in CodeXGLUE code refinement and showing notable performance on several datasets, including Defects4J. We answered several critical research questions, offering in-depth insights, discussions, and conclusions through meticulous experimentation. Our study not only advanced the current state-of-the-art in APR but has also paved the way for future research to explore newer dimensions of LLMs in software engineering. As the potential of LLMs continues to unfold, we anticipate Toggle’s methodology to inspire subsequent models, fostering even more robust, reliable, and efficient automated software engineering solutions.

### v3

@@ 3.1.2. Baseline
For our baselines, we selected the two best-performing models, based on our current knowledge. Given that the Tufano Small and Medium datasets are part of the CodeXGLUE benchmark (Leaderboard, 2023), our selection was based on the highest-ranking models from the leaderboard. For the CodeReviewer dataset, our baselines include the CodeT5 model as reported in (Li et al., 2022) and NSEdit (Hu et al., 2022) model. These models achieve the top two accuracies on the CodeReviewer dataset.
The NSEdit method (Hu et al., 2022), a 12-layer encoder-decoder with 223M parameters, expands to around 500M including the re-ranking model. It achieved 23.86% EM accuracy on the Tufano Small dataset and 13.46% on the Tufano Medium. Another baseline, CoText (Phan et al., 2021), also has 223M parameters but supports up to 11B, achieving 22.64% on the small and 15.36% on the medium dataset.
On the CodeReviewer with comment dataset, the CodeT5 model achieved an accuracy of 24.42%, as reported in the original paper (Li et al., 2022) [^1] In the CodeReviewer paper, specialized additional pre-training increased accuracy to 30.32%. Since the pre-training was custom and the dataset is not public, we compare our results with their reported CodeT5 results.. For the CodeReviewer without comment dataset, the original paper did not report any accuracy (Li et al., 2022). However, replicating their method on this dataset is straightforward and yields an accuracy of 11.48%. For both CodeReviewer with and without comment datasets, we run the NSEdit model (Hu et al., 2022) and report the accuracies in Table 1.
@@ 3.2.1. Experimental Setup
Dataset. Defects4J is a benchmark dataset consisting of real-world Java bugs from 17 projects. This dataset is widely used for evaluating automated program repair methods. From Defects4J, we primarily focus on ‘single-hunk’ bugs, which can be fixed with changes in a single contiguous block of code. This type of bug is well-suited for APR methods and has been extensively used in prior research (Lutellier et al., 2020; Jiang et al., 2021; Li et al., 2020; Ye et al., 2022; Jiang et al., 2023b). This study considers a total of 240 single-hunk bugs from the Defects4J benchmark.
Patch generation with fine-tuned model. For location prediction, we employ our fine-tuned CodeT5 bug localization model. As we did not fine-tune the model on the Defects4J dataset, to ensure diversity of sample distribution seen by the model, we fine-tune it on different datasets: Tufano Small, Tufano Medium, and GitHub (Zhu et al., 2021). For bug fixing task, we select CodeParrot, a 110M parameter model, due to its superior performance over the similarly sized CodeGPT-110M and comparable performance to larger models like CodeGen and PolyCoder. Considering the resource-intensive nature of larger models, we used the CodeParrot model fine-tuned on same three datasets.
To generate patches for Defects4J bugs, we first predict bug locations with three fine-tuned CodeT5 models. Then, similar to RQ1, we construct prompt 3 (¡truncated bug¿ [sep] ¡shared prefix¿), as shown in Figure 5. Each fine-tuned bug fixing model then generates 70 patches per buggy input sample, resulting in a total of 210 patches for each bug. This quantity is relatively small compared to the number of patches generated by most existing APR techniques (Lutellier et al., 2020; Jiang et al., 2021; Ye et al., 2022; Xia and Zhang, 2022; Jiang et al., 2023b; Meng et al., 2023) evaluated on Defects4J.
Accuracy computation via test execution. Unlike RQ1, where accuracy was determined via token-by-token exact match metric, this experiment utilizes the test cases from the Defects4J framework to validate the generated fixes. A patch is considered correct if it successfully passes all test cases. To further enhance our confidence, we manually executed all corrective patches to ensure they indeed pass all tests and effectively fix the bugs.
@@ 3.2.2. Results
CURE (Jiang et al., 2021) 18 32 37 49 52 70
RewardRepair (Ye et al., 2022) 28 39 49 58 73 85
Recoder (Zhu et al., 2021) 36 51 57 64 85 -
KNOD (Jiang et al., 2023b) 33 49 62 70 84 100
Tare (Zhu et al., 2023) - - - - - 109
AlphaRepair (Xia and Zhang, 2022) - - - - - 110
TENURE (Meng et al., 2023) - - - - - 129
@@ 3.3.1. Experimental Setup
To solely assess the impact of our four designed prompts (shown in Figure 5) on the bug fixing accuracy, we utilize the ground truth locations of buggy tokens instead of predicted ones. This method ensures a controlled setup, specifically designed to prevent location prediction errors from influencing the accuracy of bug fixing. We fine-tune six bug-fixing LLMs with each of the four prompts on the Tufano Small dataset (Tufano et al., 2019) from the CodeXGLUE benchmark (Lu et al., 2021). Finally, we report their test accuracy in Table 4.
@@ 3.3.2. Results
In Table  4, each row shows the bug-fixing accuracy of a specific LLM, and each column shows the accuracy achieved with each prompt formats. CodeGPT starts with a lowest accuracy of 16.07% using prompt 1 but shows a consistent and significant improvement across prompts, reaching 56.98% with prompt 4. CodeParrot begins at 24.5% accuracy with prompt 1, marking the highest accuracy among all models for this prompt. Accuracy for prompts 2 and 3 remains similar; however, with prompt 4, CodeParrot reaches 63.83%, the highest across all models for this prompt. CodeGen-350M exhibits a similar pattern, beginning at 22% and consistently improving to achieve up to 60.22% accuracy. PolyCoder also shows this trend, starting at 23.66% and consistently improving to achieve an accuracy of 62.29%.
@@ 4.1. Deep Learning-Based Automated Program Repair
Research in deep learning-based program repair has primarily focused on three approaches: neural machine translation (NMT)-based, edit-based, and the more recent large language model (LLM)-based methods. NMT-based approaches treat the correction of code as a translation task, where the ‘fixed’ code is translation target. This method is relatively simple to implement because both the input (buggy code) and the target (fixed code) are sequences of code, enabling direct modeling as a sequence-to-sequence (Seq2Seq) task (Chen et al., 2019; Lutellier et al., 2020; Jiang et al., 2021; Ye et al., 2022). However, a potential downside of directly predicting the fixed code is that it may encourage the models to learn the copying behavior due to the typical large overlap between the buggy and fixed code. This copying behavior causes over-fitting and overlooks the goal of editing to change the code and fix the bug.
Editing-based approaches try to directly target the portion of the code that contains bug and usually contain an explicit mechanism for bug localization. Although edits are often shorter than the fixed code, most of the existing approaches often require multiple stages of editing with complicated models (Hashimoto et al., 2018; Zhu et al., 2021; Yin et al., 2019; Tarlow et al., 2020; Jiang et al., 2023b; Zhu et al., 2023). They may also rely on a graph representation of the buggy code (Yao et al., 2021; Dinella et al., 2020), with the implicit assumption that the buggy code can be parsed into a graph, which may not be true if bugs contain syntactical errors.
Recent studies have explored large language models (LLMs) for automated program repair (APR), showing that LLMs, either as-is or with straightforward fine-tuning, have good fixing capabilities (Jiang et al., 2023a; Xia et al., 2023). However, these studies did not address bug localization or its integration with bug fixing, nor did they conduct a comprehensive study on the design of prompts for fine-tuning LLMs.
@@ 4.2. Large Language Models For Software Engineering
LLMs have been applied to several software engineering tasks in addition to APR, including code generation (Chen et al., 2022; Liu et al., 2023; Chen et al., 2023; Le et al., 2022), vulnerability detection and fixing (Wu et al., 2023; Steenhoek et al., 2023; Christou et al., 2023), fuzzing (Deng et al., 2023) and software specification generation (Xie et al., 2023). Such adoption of LLMs typically requires fine-tuning if a decent amount of data is available, or prompt design if the prompt engineering approach is utilized.
@@ 5. Threats to Validity
In our research, we have investigated the effectiveness of our bug fixing framework using several open-source, widely recognized datasets such as CodeXGLUE  (Lu et al., 2021) and CodeReviewer  (Li et al., 2022). These datasets present a diverse set of bugs collected from GitHub and span various programming languages. However, there’s a possibility that our results may not generalize across other datasets. To address this threat and ensure broader applicability, we further evaluated the performance of our fine-tuned model on the unseen dataset, Defects4J, which is a widely used real-world bug dataset.
In our large-scale experiments, we have developed various tools and scripts to facilitate our experiments, and it is conceivable that they might contain bugs. To mitigate this threat, we have used widely used PyTorch and Hugging Face libraries. Additionally, we have carried out thorough validity checks and repeated each experiment several times to confirm consistency.
In our study, we measured bug fixing accuracy using the ‘exact match’ metric, widely adopted by the research community. Specifically for the Defects4J dataset, we determined accuracy through test validation. This involved executing the test cases on the generated patches; a patch was deemed correct if it passed all test cases. We are confident that this approach provides a reasonable and reliable accuracy metric.
@@ 6. Conclusion
In this study, we have explored the intricacies of LLMs in the realm of automated program repair. To sum up, this paper makes the following contributions: (1) Granularity Shift: We introduced a new direction that localizes and fixes bugs at token granularity instead of the traditional line granularity. This innovation drastically reduces the redundancy for bug-fixing models to produce shared prefixes or suffixes. (2) Prompting Strategy: We presented a novel design of four diverse prompts to optimize LLMs as bug-fixing models. This approach proves that with the right prompt design, token-granulated localization brings a potent inductive bias, leading to substantial improvements in bug-fixing accuracy. (3) Harmonizing Discrepancies: We addressed the unique challenge posed by inconsistencies between bug localization and fixing models during LLM fine-tuning. To reconcile this, we proposed an intermediate adjustment module, further strengthening the capability of our bug-fixing models. (4) In-depth Exploration: Our extensive research into LLMs’ bug-fixing capabilities encompassed experiments with four different prompts across six LLMs and five benchmarks. The culmination of this research resulted in the creation of Toggle, our Token-Granulated Bug Localization and Repair framework. Toggle has surpassed existing state-of-the-art techniques across several benchmark assessments.
Our proposed framework, Toggle demonstrates the success of our strategy, establishing a new benchmark in CodeXGLUE code refinement and showing notable performance on several datasets, including Defects4J. We answered several research questions, offering in-depth insights, discussions, and conclusions through meticulous experimentation. Our study not only advanced the current state-of-the-art in APR but has also paved the way for future research to explore newer dimensions of LLMs in software engineering. As the potential of LLMs continues to unfold, we anticipate Toggle’s methodology to inspire subsequent models, fostering even more robust, reliable, and efficient automated software engineering solutions.